I'm not sure if this is a typo or what.. I'm really stumped. Could you help me? thanks!
sin 25degrees8'8"

Answer by ewatrrr(24785)

 
Hi,
Not a Typo, 1' = 1/60degrees and 1" = 1/3600 degrees (60" in 1')
Read 25 degrees, 8 minutes and 8 seconds
sin(25° 8'8") = sin( 25 + 8/60 + 8/3600) = sin(25.1256) to 4 decimal places